OpenStudy (anonymous):

The price of a CD is $16. If the sales tax is 4% what will be the total price of the CD?

OpenStudy (anonymous):

Are you kidding?

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

Nevermind that, anyways is this for homework/quiz/test?

OpenStudy (anonymous):

the taxi =4%= (4/100)X 16= 0.64;then the price = 16+0.64=16.64

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

16.64 16x1.04=16.64